OpenStudy (anonymous):

which of the following is a point on the line which passes through (-3,-17) and (12,-7)? A. (75,33) B. (75,35) C. (75,37) D. (75,39) E. (75,41)

OpenStudy (anonymous):

start by finding an equation of the line. slope is (-7+17)/(12+3)=10/15=2/3 using point slope form y+17=(2/3)(x+3) y=2/3x-15 plug in x=75 y=(2/3)*75-15 y=50-15=35 so (70,35) is on the line